U.V. light of wavelength 800 Å & 700 Å falls on hydrogen atoms in their ground state & liberates
electrons with kinetic energy 1.8 eV and 4 eV respectively. Calculate planck’s constant.

Grade:11

1 Answers

erra akhil
8 years ago
Dear Lord Stifner,
Given,Wavelength(W1) = 800A = 800*10-10m
Wavelength(W2) = 700A = 700*10-10m
KE1=1.8 eV = 1.8*10-19J
KE2=4 eV = 4*10-19J
We know in general,
Energy=Work Function+Kinetic Energy
here, light from uv rays with different wavelength have different energy.
Energy=hc/w.
h--planks constant
c--sped of light=3*108m/s
w--wavelength
Using the formula,
Energy=Work Function+Kinetic Energy
we get 2 eqns,on solving them we get the value of planks constant.